OpenStudy (anonymous):

Find the domain of the function h(x)=5-3x over 2x^2-14x=20

OpenStudy (lgbasallote):

\[h(x) = \frac{5-3x}{2x^2 - 14x + 20}\] this is thje question right?

OpenStudy (lgbasallote):

the*

OpenStudy (anonymous):

or this:\[h(x)=\frac{5-3x}{2x^2-14}=20\]

OpenStudy (lgbasallote):

to get the domain RESTRICTION jsut get the value of x in \[2x^2 - 14x + 20 = 0\]

OpenStudy (lgbasallote):

do you know how to find x?

OpenStudy (lgbasallote):

to find x first divide everything by 2 x^2 - 7x + 10 = 0 now can you factor that out?